OpenStudy (mjmk):

Two of the top pages when googled show different rules on exponential graphs. I'm pretty sure which is right, what do you think? -2^0 is -1 or 1? Don't rely on your calculator to get it right either. So do negative bases alternate between positive and negative outcomes? Purplemath would disagree.

OpenStudy (saifoo.khan):

-1

OpenStudy (anonymous):

-1

OpenStudy (katrinakaif):

-1 Because : -2^0 = -(2^0) = -(1)

OpenStudy (mjmk):

http://www.homeschoolmath.net/teaching/negative_zero_exponents.php then check this out

OpenStudy (katrinakaif):

Interesting..though I have seen the site before..I still would stick with -1 though.. its a "valid" argument?

OpenStudy (mjmk):

-2^2=4

OpenStudy (mjmk):

A graphics calculator will draw all negative values but then again accordng to it -1^2=-1 and we know that's nonesense

OpenStudy (mjmk):

Ahhhh the brachets help try (-2)^X alternating

OpenStudy (mjmk):

Is it about correct notation then?

OpenStudy (anonymous):

lets say a^n. For a negative if n is even you have a positive outcome. For a odd you get a negative outcome

OpenStudy (mjmk):

Yes my point but look at purple math http://www.purplemath.com/modules/graphexp4.htm says that the output is always negative. This contradicts the id that anything to power zero is 1

OpenStudy (anonymous):

if you had (-2)^0 it would be 1. Now it is like having (-1)(2)^0

OpenStudy (mjmk):

I'll be back later. Thanks for your input